This year, your brother Jack will be 2 years from being twice as old as your sister Jen. The sum of Jack's age and three times Jen's age is 66. How old is Jen?
J=Jack
E=Jen
J+2=2E
J+3E=66
J=2E-2
2E-2+3E=66
5E=66
E=13.6 This doesn't seem correct?
Thanks! :)

Answer by Alan3354(69443)

This year, your brother Jack will be 2 years from being twice as old as your sister Jen. The sum of Jack's age and three times Jen's age is 66. How old is Jen?
J=Jack
E=Jen
J+2=2E
J+3E=66
J=2E-2
2E-2+3E=66
5E=66 *** 68
E=13.6 This doesn't seem correct?
Thanks! :)
--------
It is correct, it all checks out. It's unusual to have non-integers for age problems, maybe a typo in the book or in transcribing the problem.